# Short history of signed exceptional sequences

### v1

@@ Abstract.
Whereas exceptional sequences have a long history with many well-known connections to combinatorics, signed exceptional sequences are relatively recent. The authors introduced this concept in 2017 [20], although it was retroactively realized that the category of noncrossing partitions [25] is a special case of this construction. Buan and Marsh [5] have introduced the concept of $\tau$-exceptional sequences to generalize the definitions and theorems to all finite dimensional algebras. This short paper is the story of the original concept of signed exceptional sequences for hereditary algebras and how it developed out of the two authors’ study of algebraic K-theory, link invariants, and cluster combinatorics.
@@ Introduction
Every exceptional sequence $(E_{1},\cdots,E_{k})$ produces a “perpendicular category” or, more precisely, a “Hom-Ext perpendicular category” $(E_{1}\oplus\cdots\oplus E_{k})^{\perp_{01}}$ which it the full subcategory of all modules $M$ so that $\operatorname{Hom}_{\Lambda}(E_{i},M)=0=\operatorname{Ext}^{1}_{\Lambda}(E_{i},M)$ for all $i\leq k$. Call this category ${\mathcal{W}}_{k}$. Then ${\mathcal{W}}_{k}$ is a finitely generated “wide subcategory” of $mod\text{-}\Lambda$ which is equivalent to the module category of a hereditary algebra of rank nkn-k. The exceptional sequence $E_{1},\cdots,E_{k}$ can be completed on the left by adjoining a complete exceptional sequence of ${\mathcal{W}}_{k}$. We get a complete exceptional sequence
@@
Signed exceptional sequences arose from the study of “pictures”, “picture groups” and “picture spaces”. The broad outline is as follows. In the 1970’s, for his PhD thesis work, the first author developed the concept of a “picture” and used it to define algebraic K-theory invariants for diffeomorphisms of smooth manifolds. Later, working with Kent Orr, pictures were used to settle a conjecture about link invariants. Around the turn of the century, Fomin and Zelevinsky developed the theory of cluster algebras and the second author, together with Buan, Marsh, Reineke and Reiten related this to tilting theory. They “categorified” cluster theory by creating the cluster category. Fomin-Zelevinsky clusters were shown to correspond to “cluster-tilting objects” in the cluster category. These were extensions of “tilting objects” which were sets of indecomposable modules which formed the vertices of a triangulation of the $n$-simplex. By adding shifted projective objects, this triangulation was extended to a triangulation of the $n$-sphere. The first author saw the resemblance with his “pictures” for the Steinberg group. Jerzy Weyman recognized these pictures as variants of the semi-invariant pictures from his work with Harm Derksen. The four authors: Igusa, Orr, Todorov, Weyman combined their ideas in [15], [16].
@@
There are several proofs that the picture space is a $K(\pi,1)$. The most satisfying proof shows that the space is a cubical complex which is locally CAT(0) and, therefore, a $K(\pi,1)$. To covert a simplicial complex into a cubical complex, we take the category of simplices with inclusion maps. This is the cluster morphism category! Sequences of minimal simplex inclusions are signed exceptional sequences! Every simplex (corresponding to a cluster tilting object) produced a cube of morphisms. There are $n!$ edge paths for each cube and therefore we obtain $n!$ signed exceptional sequences for every cluster-tilting object!
Outline of this paper: In Section 1, we define “pictures” and discuss their relation to homology of groups. Pictures are planar diagrams labeled with generators and relations of a group $G$. Pictures represent elements of $H_{3}G$. For the Steinberg group of a ring $R$, we have $H_{3}(St(R))\cong K_{3}(R)$. In Section 2, we discuss cluster categories and cluster-tilting objects in a cluster category. We indicate how, for finite type, this gives a triangulation of the unit sphere $S^{n-1}$ in ${\mathbb{R}}^{n}$. In Section 3, we show how the “walls” of the picture are given as domains of definition of semi-invariants on the representation space. In Section 4 we define picture groups and outline the construction of picture spaces for Dynkin quivers. One example had already been constructed by Loday. He constructed what we now call the picture space of type $A_{n}$ by identifying faces of the Stasheff associahedron. We go over Loday’s example and generalize it. We also present our calculation of the homology of picture groups of type $A_{n}$ assuming the picture space is a $K(\pi,1)$ for the picture group. In Section 5 we explain how dualizing a simplex makes it into a cube. So, picture space is seen to be a union of cubes, one for every cluster-tilting object of the category. In Section 6, we construct the cluster morphism category and show that its geometric realization is a union of cubles. Finally, in Section 7 we define signed exceptional sequences and give examples.
@@ 1. Pictures and homology of groups
Let $G$ be a group with a presentation $G=\left<{\mathcal{X}}|{\mathcal{Y}}\right$> where elements of $\mathcal{Y}$} (the relations) are cyclically reduced words in ${\mathcal{X}}\cup{\mathcal{X}}^{-1}$ and ${\mathcal{Y}}\cap{\mathcal{Y}}^{-1}=\emptyset$, A picture for $G$ is a planar diagram consisting of vertices labeled with elements of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$ together with a “base point direction” and oriented edges labeled with elements of $\mathcal{X}$} so that the labels of the edges incident to any vertex, read counterclockwise from the base point direction given an element of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$.
@@
There are three oriented edges labeled $x$. The labels of the vertices are elements of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$ given by reading the incoming edge labels (and inverse outgoing labels) starting at the base point direction indicated with $\ast$. The base point direction is omitted when it is uniquely determined. We are using the drawing convention that edges should be curved to the right.
@@
instead of the usual $xyx^{-1}y^{-1}$. In [26] we took our old paper and redrew the pictures using the representation theoretic convention to convert the original Morse theory pictures into semi-invariant pictures. Here is a picture for $G={\mathbb{Z}}^{3}$.
$\ast\astxxyyzz$
We use the convention that smooth curves have the same label throughout. Also, the base point directions are uniquely determined and are omitted. We left two inverse relations $[x,y]$ and $x,y]^{-1}$. The other four vertices are $[y,z],[z,x]$ and their inverses.
@@
A picture $P$ for $G=\left<{\mathcal{X}}|{\mathcal{Y}}\right$> represents an element of $H_{3}(G)$, the third homology of $G$ if, for every $\in{\mathcal{Y}}, yy$ and $y^{-1}$ occur the same number of times. Furthermore, every element of $H_{3}(G)$ is represented by such a picture.
@@
Combining the last two theorem we conclude that elements of $K_{3}({\mathbb{Z}}\pi)$ were represented by certain pictures for the Steinberg group of ${\mathbb{Z}}\pi$. But this is not good enough since 2-parameter Morse theory produces arbitrary pictures for the Steinberg group. We needed the following.
@@
There are five (series of) second order relation. Two of them are shown in Figure 1. We continue using the convention that arcs are oriented to curve right and base point directions are uniquely determined.
@@ 2. Cluster categories and cluster tilting objects
The general result it the following.
@@
For any hereditary algebra $\Lambda$ of finite representation type having $n$ simple modules, there is a triangulation of the unit sphere $S^{n-1}$ with vertices corresponding to the indecomposable objects of the cluster category and so that $n$ vertices span a top dimensional simplex in this triangulation if they form a cluster tilting object, i.e., if the vertices are pairwise Ext-orthogonal.
@@ 3. Semi-invariant pictures
The weight of $\chi_{ij}$ is defined to be $(i,j)\in{\mathbb{Z}}^{2}$. So, the weight of $\det:\operatorname{Hom}_{K}(K^{n},K^{n})\to K$ is (1,1)(1,-1). (See [16, Appendix B] for discussion of nondeterminantal weights.)
@@
Take a $KQ$-module $V$ considered as representation of $Q:1\leftarrow 2\leftarrow 3$:
@@
The domains of these semi-invariants (the set of dimension vectors of $V$ on which the semi-invariants are nonzero) are indicated in Figure 3. Conversion to “projective” semi-invariants goes like this.
@@
Figure 4. This is the $g$-vector fan for $A_{3}$ in ${\mathbb{R}}^{3}$ interested with the unit sphere $S^{2}$ and stereographically projected to the plane. Thus, hyperplane, such as $D(S_{1})$ become circles in this figure. The half-plane $D(P_{2})$ becomes a semi-circle. This figure is homeomorphic to the figure in Example 2.2. We call this the semi-invariant picture for $A_{3}$. This is also the representation theoretic version of the “second order Steinberg relation” shown in Figure 2.
@@ 4. Picture groups and picture spaces
Given a picture $P$ for some group $G=\left<{\mathcal{X}}\,|{\mathcal{Y}}\right$>, the picture group of $P$ is the group $G(P)=\left<{\mathcal{X}}_{0}\,{\mathcal{Y}}_{0}\right$> where ${\mathcal{X}}_{0}\subset{\mathcal{X}}$ is the set of edge labels which actually occurs in the picture $P$ and ${\mathcal{Y}}_{0}\subset{\mathcal{Y}}$ is the set of relations which label the vertices of $P$.
@@ 4.1. Loday’s construction
For example, $Sta_{3}$ has only three generators $x_{12},x_{23}$ and $x_{13}$ with only one relation $x_{13}=[x_{12},x_{23}]$. So, $Sta_{3}$ is the free group on two generators. Loday obtained his Stasheff group by putting labels on edges and faces of the Stasheff associahedron and collapsing all vertices to one vertex, identifying edges with the same labels, taking faces to be relations and identifying faces with the same relations. For example, $K_{3}$ is a pentagon with oriented edges labeled $12,23,13$ as shown in Figure 6.
@@
Figure 6. On the left, Loday composes arrow right to left to get the relation $(12)(23)=(23)(13)(12)$ which corresponds to the relation $x_{12}x_{23}=x_{23}x_{13}x_{12}$ which is an equality of row operations corresponding to the dual hand slide picture on the right.
@@ 4.2. Picture group for Dynkin quivers
The edges of the dual of $\Sigma(Q)$ are the lines that go from the barycenters of two n1n-1 dimensional simplex which shares an n2n-2 dimensional face. Recall from Theorem 3.2 that the n2n-2 dimensional simplices form the domains of semi-invariants $D(M)$. The edge goes through the barycenter of that face. See Figure 9. We use the notation $M_{\alpha}$ to denote the indecomposable module with dimension vector $\alpha=\operatorname{\underline{dim}}M_{\alpha}$. It is well-know [12] that these vectors form the set of positive roots of the root system with Dynkin diagram the underlying graph of the valued quiver $Q$ of $\Lambda$.
@@
A dual 2-cell $E(\rho)$ of $D\Sigma(Q)$ is a $k$-sided polygons centered at the barycenter of a codimension-2 simplex $\rho$ in $\Sigma(Q)$ where $k=4,5,6$, or $8$. Figure 10 below illustrates the case $k=4$ and Figure 8 shows the case $k=6$ with $\rho$ being the center point. Looking at Figure 10 (and Figure 7) we see that the parallel sides of a dual square have the same labels. We also see that, when vertices are identified and edges with the same label are identified, any dual square will become a torus. Also, if two torus’s have the same label, they are identified. For example, in Figure 7 two squares are labeled with $12,34$. These are pasted together but with the opposite orientation.
@@
We need one more thing: Which pairs $\alpha,\beta$ form these dual squares? We go back to Loday’s construction in order to generalize it. Let $Q$ be the quiver of type $A_{n}$ shown in (4.1). The path algebra $KQ$ has projective modules $P_{j}$ given by $K$ at each vertex $i\leq j$ and 0 elsewhere. So, the support of $P_{j}$ is $(0,j]$. Then, all indecomposable modules are given by $M_{ij}=P_{j}/P_{i}$ where $P_{0}:=0$. This has support $(i,j]$. In particular, $P_{j}=M_{0j}$ and the simple module $S_{j}$ is $M_{j-1,j}$. Note that $M_{ij}$ is a submodule of $M_{k\ell}$ if and only $i=k$ and $j\leq\ell$. Similarly $M_{ij}$ is a quotient of $M_{k\ell}$ if and only if $j=\ell$ an $k\leq i$. This implies that $M_{ij},M_{k\ell}$ are Hom-orthogonal, i.e.,
@@
The Stasheff group $Sta_{n}$ is isomorphic to the picture group of $A_{n-1}$ with straight orientation.

### v2

@@ Abstract.
Whereas exceptional sequences have a long history with many well-known connections to combinatorics, signed exceptional sequences are relatively recent. The authors introduced this concept in 2017 [22] although it was retroactively realized that the category of noncrossing partitions [27] is a special case of this construction. Buan and Marsh [6] have introduced the concept of $\tau$-exceptional sequences to generalize the definitions and theorems to all finite dimensional algebras. This short paper is the story of the original concept of signed exceptional sequences for hereditary algebras and how it developed out of the two authors’ study of algebraic K-theory, link invariants and cluster combinatorics.
@@ Introduction
Every exceptional sequence $(E_{1},\cdots,E_{k})$ produces a “perpendicular category” or, more precisely, a “Hom-Ext perpendicular category” $(E_{1}\oplus\cdots\oplus E_{k})^{\perp_{01}}$ which is the full subcategory of all modules $M$ so that $\operatorname{Hom}_{\Lambda}(E_{i},M)=0=\operatorname{Ext}^{1}_{\Lambda}(E_{i},M)$ for all $i\leq k$. Call this category ${\mathcal{W}}_{k}$. Then ${\mathcal{W}}_{k}$ is a finitely generated “wide subcategory” of $mod\text{-}\Lambda$ [30] which is equivalent to the module category of a hereditary algebra of rank nkn-k. The exceptional sequence $E_{1},\cdots,E_{k}$ can be completed on the left by adjoining a complete exceptional sequence of ${\mathcal{W}}_{k}$. We get a complete exceptional sequence
@@
Signed exceptional sequences arose from the study of “pictures”, “picture groups” and “picture spaces”. The broad outline is as follows. In the 1970’s, for his PhD thesis work, the first author developed the concept of a “picture” and used it to define algebraic K-theory invariants for diffeomorphisms of smooth manifolds. Later, working with Kent Orr, pictures were used to settle a conjecture about link invariants. Around the turn of the century, Fomin and Zelevinsky developed the theory of cluster algebras [11, 12, 13] and the second author, together with Buan, Marsh, Reineke and Reiten related this to tilting theory. They “categorified” cluster theory by creating the cluster category [5]. This was also done in the $A_{n}$ case by Caldero, Chapoton and Schiffler [7]. Fomin-Zelevinsky clusters were shown to correspond to “cluster-tilting objects” in the cluster category. These were extensions of “tilting objects” which were sets of indecomposable modules which formed the vertices of a triangulation of the $n$-simplex. By adding shifted projective objects, this triangulation was extended to a triangulation of the $n$-sphere. The first author saw the resemblance with his “pictures” for the Steinberg group. Jerzy Weyman recognized these pictures as variants of the semi-invariant pictures from his work with Harm Derksen. The four authors: Igusa, Orr, Todorov, Weyman combined their ideas in [17], [18].
@@
There are several proofs that the picture space is a $K(\pi,1)$. The most satisfying proof shows that the space is a cubical complex which is locally CAT(0) and, therefore, a $K(\pi,1)$. To covert a simplicial complex into a cubical complex, we take the category of simplices with inclusion maps. This is the cluster morphism category. Sequences of minimal simplex inclusions are signed exceptional sequences. Every simplex (corresponding to a cluster tilting object) produced a cube of morphisms. There are $n!$ edge paths for each cube and therefore we obtain $n!$ signed exceptional sequences for every cluster-tilting object.
Outline of this paper: In Section 1, we define “pictures” and discuss their relation to homology of groups. Pictures are planar diagrams labeled with generators and relations of a group $G$. Pictures represent elements of $H_{3}G$, the third integral homology of the group $G$ (Theorem 1.4). For the Steinberg group of a ring $R$, $H_{3}(St(R))$ is isomorphic to $K_{3}(R)$, the third algebraic $K$-theory group of $R$ (Theorem 1.5). In Section 2, we discuss cluster categories and cluster-tilting objects in a cluster category. We indicate how, for finite type, this gives a triangulation of the unit sphere $S^{n-1}$ in ${\mathbb{R}}^{n}$. In Section 3, we show how the “walls” of the picture are given as domains of definition of semi-invariants on the representation space. In Section 4 we define picture groups and outline the construction of picture spaces for Dynkin quivers. One example had already been constructed by Loday. He constructed what we now call the picture space of type $A_{n}$ by identifying faces of the Stasheff associahedron. We go over Loday’s example and generalize it. We also present our calculation of the homology of picture groups of type $A_{n}$ assuming the picture space is a $K(\pi,1)$ for the picture group. In Section 5 we explain how dualizing a simplex makes it into a cube. So, picture space is seen to be a union of cubes, one for every cluster-tilting object of the category. In Section 6, we construct the cluster morphism category and show that its geometric realization is a union of cubles. Finally, in Section 7 we define signed exceptional sequences and give examples.
@@ 1. Pictures and homology of groups
Let $G$ be a group with a presentation $G=\left<{\mathcal{X}}|{\mathcal{Y}}\right$> where elements of $\mathcal{Y}$} (the relations) are cyclically reduced words in ${\mathcal{X}}\cup{\mathcal{X}}^{-1}$ and ${\mathcal{Y}}\cap{\mathcal{Y}}^{-1}=\emptyset$, A picture for $G$ is a planar multigraph consisting of labeled vertices and labeled edges with additional structure. Each vertex is labeled with an element of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$ and decorated with an asterisk $\ast$ in one chosen region between incident edges. We call this the “base point direction” at the vertex. In some cases we decorate a vertex with its vertex label which we place in the region of the base point direction and, in that case, the asterisk becomes redundant and we delete it. Each edge is oriented and labeled with an element of $\mathcal{X}$}. At each vertex we require that the labels of the edges incident to that vertex, with the label inverted when the edge points away from the vertex, read counterclockwise from the base point direction at the vertex, give a word in ${\mathcal{X}}\cup{\mathcal{X}}^{-1}$ which is the element of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$ which is the label at that vertex.
We use the drawing convention that indicates the orientation of edges by having them curve to the right. For example, circles should be oriented clockwise.
In some cases, such as Example 1.3 below, the edge labels will uniquely determine the vertex labels together with their base point directions in the sense that there is only one choice of these vertex structures which is consistent with the definition. By this we mean there is a unique region next to the vertex so that, reading the edge labels counterclockwise around the vertex starting in this region, we obtain an element of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$. That unique region is the base point direction and the element of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$ that we obtain is the vertex label. In that case, we often omit these uniquely determined labels. In Example 1.3, we could have deleted the vertex labels since they are uniquely determined. However, for better understanding of the basic definitions, we have inserted the vertex labels and placed them in the base point direction of each vertex.
@@
There are three oriented edges labeled $x$. The labels of the vertices are elements of ${\mathcal{Y}}\cup{\mathcal{Y}}^{-1}$ given by reading the incoming edge labels (and inverse outgoing labels) counterclockwise starting at the base point direction indicated with $\ast$. In this picture, there are three possible base point directions at each vertex. They give the same relation ($x^{3}$ on the right and $x^{-3}$ on the left). However, if we were to move the base point direction at either vertex, the picture would become a different picture not isomorphic to the given picture. The orientation and curvature of edges agrees with the drawing convention that edges should be curved to the right.
@@
instead of the usual $xyx^{-1}y^{-1}$. In [28] we took our old paper and redrew the pictures using the representation theoretic convention to convert the original Morse theory pictures into semi-invariant pictures. Here is a picture for $G={\mathbb{Z}}^{3}$ with vertices and edges labeled.
[…]
All six vertices are labeled with relations or inverse relations. The labels of the vertices are placed in the base-point direction. For example, at the lower right vertex, we have the label $b=[y,z]$. The location of the base-point direction is uniquely determined in this case since the four possible words that can be read counterclockwise at that vertex are:
$yzy^{-1}z^{-1},zy^{-1}z^{-1}y,y^{-1}z^{-1}yz,z^{-1}yzy^{-1}.$
Only the last one $y,z]:=z^{-1}yzy^{-1}$ is in our list of relation or inverse relations. So, $b=[y,z]$ is the relation at that vertex and the word order dictates that we start at the location (space between edges) marked “bb”.
We use the convention that smooth curves have the same label throughout. Also, when the base point directions are uniquely determined and the relations at the vertices are uniquely determined as in this example (but not in Example 1.2), we can omit them from the notation. In this picture, the vertex labels $a,b,c$ and their inverses can be safely omitted.
@@
A picture $P$ for $G=\left<{\mathcal{X}}|{\mathcal{Y}}\right$> represents an element of $H_{3}(G)$, the third homology of $G$ if, for every $\in{\mathcal{Y}}, yy$ and $y^{-1}$ occur the same number of times as vertex labels. Furthermore, every element of $H_{3}(G)$ is represented by such a picture.
@@
Combining the last two theorems we conclude that elements of $K_{3}({\mathbb{Z}}\pi)$ were represented by certain pictures for the Steinberg group of ${\mathbb{Z}}\pi$. But this is not good enough since 2-parameter Morse theory produces arbitrary pictures for the Steinberg group. We needed the following.
@@
There are five (series of) second order relations. Two of them are shown in Figure 1. We continue using the convention that arcs are oriented to curve right and base point directions are uniquely determined.
@@ 2. Cluster categories and cluster tilting objects
The general result is the following.
@@
For any hereditary algebra $\Lambda$ of finite representation type having $n$ simple modules, there is a triangulation of the unit sphere $S^{n-1}$ with vertices corresponding to the indecomposable objects of the cluster category and so that $n$ vertices span a top dimensional simplex in this triangulation if and only if they form a cluster tilting object, i.e., if the vertices are pairwise Ext-orthogonal.
@@ 3. Semi-invariant pictures
The weight of $\chi_{ij}$ is defined to be $(i,j)\in{\mathbb{Z}}^{2}$. So, the weight of $\det:\operatorname{Hom}_{K}(K^{n},K^{n})\to K$ is (1,1)(1,-1). (See [18, Appendix B] for discussion of nondeterminantal weights.)
@@
Take a $KQ$-module $V$ considered as a representation of the quiver $Q:1\leftarrow 2\leftarrow 3$:
@@
These semi-invariants are not always defined. The set of dimension vectors of $V$ for which these semi-invariants are defined and nonzero form the integer points of a closed convex set which we call the domain of the semi-invariants. These domains are indicated in Figure 3. Conversion to “projective” semi-invariants goes like this.
@@
Figure 4. This is the $g$-vector fan [13, 37] for $A_{3}$ in ${\mathbb{R}}^{3}$ intersected with the unit sphere $S^{2}$ and stereographically projected to the plane. Thus, hyperplanes, such as $D(S_{1})$ become circles in this figure. The half-plane $D(P_{2})$ becomes a semi-circle. This figure is homeomorphic to the figure in Example 2.2. We call this the semi-invariant picture for $A_{3}$. This is also the representation theoretic version of the “second order Steinberg relation” shown in Figure 2.
@@ 4. Picture groups and picture spaces
Given a picture $P$ for some group $G=\left<{\mathcal{X}}\,|\,{\mathcal{Y}}\right$>, the picture group of $P$ is the group $G(P)=\left<{\mathcal{X}}_{0}\,|\,{\mathcal{Y}}_{0}\right$> where ${\mathcal{X}}_{0}\subset{\mathcal{X}}$ is the set of edge labels which actually occur in the picture $P$ and ${\mathcal{Y}}_{0}\subset{\mathcal{Y}}$ is the set of relations which label the vertices of $P$.
@@ 4.1. Loday’s construction
For example, $Sta_{3}$ has only three generators $x_{12},x_{23}$ and $x_{13}$ with only one relation $x_{13}=[x_{12},x_{23}]$. So, $Sta_{3}$ is the free group on two generators. Loday obtained his Stasheff group by putting labels on edges and faces of the Stasheff associahedron [40] and collapsing all vertices to one vertex, identifying edges with the same labels, taking faces to be relations and identifying faces with the same relations. For example, $K_{3}$ is a pentagon with oriented edges labeled $12,23,13$ as shown in Figure 6.
@@
Figure 6. On the left, Loday composes arrows right to left to get the relation $(12)(23)=(23)(13)(12)$ which corresponds to the relation $x_{12}x_{23}=x_{23}x_{13}x_{12}$ which is an equality of row operations corresponding to the dual hand slide picture on the right.
@@ 4.2. Picture group for Dynkin quivers
The edges of the dual of $\Sigma(Q)$ are the lines that go from the barycenters of two n1n-1 dimensional simplices which share an n2n-2 dimensional face. Recall from Theorem 3.2 that the n2n-2 dimensional simplices form the domains of semi-invariants $D(M)$. The edge goes through the barycenter of that face. See Figure 9. We use the notation $M_{\alpha}$ to denote the indecomposable module with dimension vector $\alpha=\operatorname{\underline{dim}}M_{\alpha}$. It is well-known [14] that these vectors form the set of positive roots of the root system with Dynkin diagram the underlying graph of the valued quiver $Q$ of $\Lambda$.
@@
A dual 2-cell $E(\rho)$ of $D\Sigma(Q)$ is a $k$-sided polygons centered at the barycenter of a codimension-2 simplex $\rho$ in $\Sigma(Q)$ where $k=4,5,6$, or $8$. Figure 10 below illustrates the case $k=4$ and Figure 8 shows the case $k=6$ with $\rho$ being the center point. Looking at Figure 10 (and Figure 7) we see that the parallel sides of a dual square have the same labels. We also see that, when vertices are identified and edges with the same label are identified, any dual square will become a torus. Also, if two toruses have the same label, they are identified. For example, in Figure 7 two squares are labeled with $12,34$. These are pasted together but with the opposite orientation.
@@
We need one more thing: Which pairs $\alpha,\beta$ form these dual squares? We go back to Loday’s construction in order to generalize it. Let $Q$ be the quiver of type $A_{n}$ shown in (4.1). The path algebra $KQ$ has projective modules $P_{j}$ given by $K$ at each vertex $i\leq j$ and 0 elsewhere. So, the support of $P_{j}$ is $(0,j]$. Then, all indecomposable modules are given by $M_{ij}=P_{j}/P_{i}$ where $P_{0}:=0$. This has support $(i,j]$. In particular, $P_{j}=M_{0j}$ and the simple module $S_{j}$ is $M_{j-1,j}$. Note that $M_{ij}$ is a submodule of $M_{k\ell}$ if and only if $i=k$ and $j\leq\ell$. Similarly $M_{ij}$ is a quotient of $M_{k\ell}$ if and only if $j=\ell$ an $k\leq i$. This implies that $M_{ij},M_{k\ell}$ are Hom-orthogonal, i.e.,
@@
The Stasheff group $Sta_{n}$ is isomorphic to the picture group of $A_{n-1}$ with linear orientation.